# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v3

@@ F.2 Dataset Construction Method
To construct our dataset of adversarial prompts, we followed a two-stage process: prompt generation using GPT-4o (OpenAI et al., 2024) and audio synthesis using F5-TTS (Chen et al., 2024). First, we leveraged OpenAI’s GPT-4o (the multimodal GPT-4 with voice capabilities) to automatically generate a diverse set of harmful or policy-violating user queries for each category of attack. GPT-4o was chosen for prompt generation due to its advanced language capabilities and alignment safeguards, ensuring that the produced queries are realistic and varied. We crafted a consistent prompt template to guide GPT-4o in creating these adversarial questions. This template provided the scenario context (enterprise, financial, medical, or illegal), specified the speaker’s gender (male or female) and emotional tone (angry or sad), and instructed GPT-4o to produce a user query attempting to break the system’s rules. The template was designed to simulate a malicious user’s perspective, instructing GPT-4o to create challenging queries that, while mimicking harmful intent, adhered to its content generation policies during the generation process. An example of the prompt construction template is shown below:
Using this template (with appropriate substitutions for each scenario and persona), we generated the full set of prompts. In total, we obtained 300 text prompts for the Jailbreak category and 300 for the Illegal Activity Guidance category. For the Jailbreak prompts, we generated 100 queries for each of the three scenarios (Enterprise, Financial, Medical). Within each scenario’s 100 prompts, we ensured an even coverage of the four gender–emotion combinations: 25 prompts each for Male-Angry, Male-Sad, Female-Angry, and Female-Sad. Similarly, for the Illegal Activity Guidance category, we produced 300 prompts evenly divided into 75 for each of the four combinations (male-angry, male-sad, female-angry, female-sad). This balancing ensures that our dataset controls for speaker gender and emotional tone, allowing us to analyze their effects on attack success. After generating the textual prompts, we converted each prompt into spoken audio using the F5-TTS (Chen et al., 2024) text-to-speech system. F5-TTS is a state-of-the-art TTS model capable of high-quality voice cloning with emotional expressiveness. For each prompt, we selected a voice profile that matched the intended gender and could convey the specified emotion (angry or sad). The text of the prompt was then synthesized into an audio file using F5-TTS, resulting in a spoken version of the query. This audio conversion is crucial because our attack evaluations are conducted via GPT-4o’s voice interface. Using F5-TTS to produce natural and emotive speech ensures the audio prompts closely mimic real user utterances in content and delivery.
@@ F.4 Evaluation Method and Metrics
Harmful Response Rate (HRR):
Here, is the number of test samples, and is the GPT-4o score for the -th response. A higher DSR reflects stronger safety alignment, while a lower HRR indicates fewer harmful completions. These two complementary metrics provide a comprehensive view of model safety under adversarial audio input. The safety of ALLMs is evaluated using these datasets and the safety metrics detailed in Section˜F.3, with results presented in Section˜F.5.
@@ F.5 Additional Result
| Male Angry | 113 | 125 | 80 | 95 | 98.0 | 122.0 | 125.0 | 46.0 | 122.0 |
| Male Sad | 89 | 117 | 77 | 103 | 85.0 | 121.0 | 125.0 | 83.0 | 109.0 |
| Female Angry | 80 | 125 | 89 | 103 | 86.0 | 125.0 | 123.0 | 64.0 | 106.0 |
| Female Sad | 89 | 119 | 95 | 80 | 84.0 | 118.0 | 124.0 | 64.0 | 103.0 |
@@
| Male Angry | 88 | 85 | 102 | 101 | 105.0 | 112.0 | 122.0 | 60.0 | 80.0 |
| Male Sad | 91 | 107 | 98 | 98 | 113.0 | 123.0 | 123.0 | 104.0 | 82.0 |
| Female Angry | 93 | 118 | 104 | 96 | 100.0 | 121.0 | 123.0 | 71.0 | 90.0 |
| Female Sad | 100 | 109 | 99 | 101 | 113.0 | 118.0 | 123.0 | 104.0 | 82.0 |
@@
| Male Angry | 106 | 118 | 108 | 113 | 109.0 | 123.0 | 119.0 | 86.0 | 101.0 |
| Male Sad | 104 | 107 | 103 | 92 | 111.0 | 113.0 | 118.0 | 96.0 | 84.0 |
| Female Angry | 90 | 117 | 95 | 94 | 110.0 | 117.0 | 124.0 | 82.0 | 92.0 |
| Female Sad | 104 | 112 | 101 | 109 | 115.0 | 118.0 | 115.0 | 112.0 | 84.0 |
@@
| Male Angry | 315 | 375 | 368 | 372 | 372.0 | 375.0 | 375.0 | 192.0 | 274.0 |
| Male Sad | 329 | 361 | 344 | 360 | 346.0 | 372.0 | 375.0 | 322.0 | 232.0 |
| Female Angry | 270 | 370 | 335 | 358 | 343.0 | 370.0 | 374.0 | 215.0 | 269.0 |
| Female Sad | 243 | 364 | 340 | 353 | 356.0 | 370.0 | 375.0 | 348.0 | 240.0 |
@@
| Male Angry | 99.2 | 125 | 124 | 125 | 125 |
| Male Sad | 97.6 | 124 | 125 | 125 | 124 |
| Female Angry | 99.2 | 123 | 125 | 125 | 125 |
| Female Sad | 100 | 123 | 125 | 125 | 125 |
@@
| Male Angry | 100 | 125 | 125 | 125 | 125 |
| Male Sad | 98.4 | 123 | 123 | 125 | 125 |
| Female Angry | 100 | 124 | 124 | 125 | 125 |
| Female Sad | 98.4 | 124 | 123 | 124 | 122 |
@@
| Male Angry | 98.4 | 123 | 125 | 125 | 125 |
| Male Sad | 96 | 121 | 123 | 124 | 124 |
| Female Angry | 99.2 | 123 | 123 | 125 | 123 |
| Female Sad | 96.8 | 124 | 122 | 123 | 125 |
@@
| Male Angry | 100 | 375 | 375 | 375 | 375 |
| Male Sad | 100 | 375 | 375 | 375 | 375 |
| Female Angry | 100 | 375 | 374 | 372 | 374 |
| Female Sad | 99.7 | 375 | 375 | 375 | 372 |
@@ G.2 Dataset Construction Method
Direct Privacy Leakage Dataset (600 samples). For the direct privacy leakage dataset,we created six categories of sensitive personal information: bank account numbers, home addresses, phone numbers, phone passwords, driver’s license numbers, and social security numbers (Wang et al., 2023a). For each category, we first prompted Grok to generate textual data containing personal names and corresponding private information (xAI, 2025). The generated data were then further randomized to enhance diversity and reduce identifiability. Based on this preliminary dataset, Grok was instructed to generate realistic dialogue scenarios in which each type of private information might naturally appear. Subsequently, Grok produced context-appropriate conversational texts for each scenario. Finally, we selected voice samples from the Common Voice dataset and used F5-TTS to synthesize the corresponding audio files (Chen et al., 2024; Ardila et al., 2020). Each category contained 100 entries, resulting in a total of 600 test samples.
@@ G.5 Experiment Design and Results
| Average | 80.4 | – | 70.7 | – | 77.4 | – | 90.3 | – | 78.7 | – | 78.5 | – | 23.6 | – | 29.5 | – | 56.9 | – |
Note: This table shows inference accuracy ONLY for samples where privacy was leaked. Higher values indicate greater privacy threat. Red
arrows () indicate higher threat than average; blue arrows () indicate lower threat. ‘’ indicates insufficient leaked samples ( 10).
@@ H.2 Dataset Construction Method
For Adversarial Robustness: Specific algorithms are applied to generate adversarial audio samples by adding imperceptible perturbations designed to mislead models.
@@ I.2 Dataset Construction Method
Voice Cloning Spoofing (200 samples). We collected 100 samples of voice wake-up words from an existing dataset (He et al., 2022), with each sample consisting of approximately 2 seconds of wake-up word audio. These samples constitute the Voice Cloning Spoofing test set. Detailed wake-up words and their corresponding developers are listed in Table˜25. Each sample was cloned using F5-TTS (Chen et al., 2024), and we constructed paired audio clips composed of the original audio (used as the legitimate user’s enrollment input) and the cloned audio (used as the attack input attempting to bypass authentication). In addition, we designed two distinct authentication scenarios for the experiments—Awake Assistant and Open The Door. In the former, the wake-up words are used to authenticate an AI voice assistant, while in the latter, they are used to activate access to a high-security laboratory door, with an emphasis on the consequences of authentication failures. Through this experimental setup, we aim to explore how different system prompt formulations affect an ALLM’s ability to defend against voice cloning spoofing under identical audio inputs.
@@ I.4 Evaluation Method and Metrics
Our evaluation across all experiments primarily relies on the following statistical metrics:
Experiment I: We measure the False Acceptance Rate (FAR), defined as the proportion of unauthorized voices incorrectly accepted by the system.
Experiment II: We calculate the authentication pass rate.
Experiment III: We compare authentication pass rates across different experimental groups.
False Acceptance Rate (FAR) – the proportion of unauthorized voices accepted by the system.
@@ I.5 Additional Result
| Ultravox | 91 | 53 | |
| Qwen2-Audio | 15 | 0 | |
| MiniCPM-o 2.6 | 27 | 14 | |
| Step-Fun | 78 | 78 | |
| Qwen2.5-omni | 100 | 100 | |
| Kimi-Audio | 92 | 59 | |
| OpenS2S | 100 | 0 | |
| Step-Audio2 | 91 | 7 | |
| Closed-source | Gemini-1.5 pro | 100 | 33 |
| GPT-4o Audio | 33 | 0 | |
| GPT-4o mini Audio | 8 | 20 | |
| Gemini-2.5 Flash | 94 | 84 | |
| Gemini-2.5-Pro | 99 | 80 |
Note: and represent the change in the proportion of failed defenses against voice cloning spoofing when comparing the ’Open the door’ scenario with stricter system prompts to the ’Awake Assistant’ scenario. Lower numbers represent better defense effectiveness.
@@
| Open-source | SALMONN | 74 | – | 93 | 92 |
| Ultravox | 5 | – | 43 | 41 | |
| Qwen2-Audio | 58 | – | 29 | 40 | |
| MiniCPM-o 2.6 | 76 | – | 57 | 44 | |
| Step-Fun | 11 | – | 3 | 2 | |
| Qwen2.5-omni | 36 | – | 36 | 64 | |
| Kimi-Audio | 21 | – | 24 | 14 | |
| OpenS2S | 3 | – | 34 | 33 | |
| Step-Audio2 | 36 | – | 85 | 80 | |
| Closed-source | Gemini-1.5 pro | 4 | 6 | 5 | 0 |
| GPT-4o Audio | 2 | 6 | 0 | 0 | |
| GPT-4o mini Audio | 0 | 9 | 0 | 0 | |
| Gemini-2.5 Flash | 3 | 4 | 7 | 19 | |
| Gemini-2.5-Pro | 5 | 0 | 3 | 0 | |
| Average | 23.9 | 5.0 | 29.9 | 30.6 | |
@@
Note: Values show false acceptance rate (FAR) (%) with performance indicators relative to average values. indicates lower FAR than average (better security performance); indicates higher FAR than average (worse security performance). Subscripts show absolute difference from average. “–” indicates the subset of the dataset was used exclusively to evaluate model performance across different modalities (audio and text). The evaluation was restricted to closed-source models to streamline the analysis.“bg” = with background audio,“no bg” = without background audio.
@@
Voice Cloning Spoofing. In Table˜26, we analyze the experimental results of all open-source and closed-source models under two scenarios with different levels of text prompt flexibility. It can be observed that most models perform better in the “Open The Door” scenario than in the “Awake Assistant” scenario. with a significant decrease in the number of samples that failed to defend against voice cloning spoofing. This is particularly evident for Ultravox, Gemini-1.5 Pro, and Gemini-2.5 Pro. This indicates that even in Audio-based Large Language Models (ALLMs) where audio is the primary input, the accuracy of text prompts still plays a significant role. Furthermore, this has implications for the downstream applications of ALLMs: for scenarios involving security, authentication, etc., designing a strict and precise prompt may lead to a considerable improvement in model performance.
@@
Identity Verification Bypass. From the Table˜27, it can be observed that closed-source models are harder to deceive compared to open-source models. Among them, GPT-4o mini Audio performs the best, with a FAR (False Acceptance Rate) of only 0%. Among all closed-source models, SALMONN performs the worst, with a FAR as high as 74%. These results indicate that even without providing complete or explicit authentication information, voice models still have a high probability of passing identity verification, which poses a significant security risk.In the Table˜27, we also investigated the FAR metric in pure text mode, which is labeled as the “Text” column. This represents using the corresponding text-based model of the audio model to perform inference on the text version of the identity verification bypass dataset. It can be observed that, in general, the FAR is higher in text mode compared to audio mode. This suggests that the additional paralinguistic information present in the speech modality, such as emotional cues or prosodic features, may contribute positively to the authentication performance of the model.
@@
Hybrid Spoofing. In this configuration, we simulate social engineering attacks combined with background audio that may occur in real authentication scenarios, aiming to study the impact of background sounds on the verification outcome. The experimental results show that the influence of added background audio on model performance does not follow an obvious pattern. For instance, Qwen2-Audio’s FAR decreases by 11%, whereas Gemini-1.5 Pro’s FAR increases instead.
@@
With the rapid increase in parameter and data scales, text-only large language models (LLMs) have achieved groundbreaking progress in language understanding and generation, as exemplified by models such as GPT‑4 and the Gemini series (Achiam et al., 2023; Team et al., 2023). Building on this, researchers explored cross-modal alignment by integrating visual information into unified representation spaces. This led to early models like CLIP (Radford et al., 2021) and Flamingo (Alayrac et al., 2022), and later, models such as GPT-4V and Gemini capable of processing high-resolution images and long contexts. Recently, ALLMs have further expanded the input modalities by incorporating temporal acoustic features (such as Mel-spectrograms, log-power spectra, or variable-length waveforms) for joint modeling with semantic tokens (Yang et al., 2024c). In contrast to the visual modality, audio signals exhibit high dynamic range and transient variations in both time and frequency domains. Consequently, most ALLMs adopt separate time-frequency encoders or discretizing acoustic tokenizers to capture rich attributes such as timbre, rhythm, and scene (Hu et al., 2024; Du et al., 2023). Representative models include Qwen2‑Audio with its pipeline-style natural language prompt pre-training (Chu et al., 2024), SALMONN with a unified "auditory-language-music" framework (Tang et al., 2024), and WavLLM with a dual-encoder plus Prompt-LoRA adaptation mechanism (Hu et al., 2024). After cross-modal alignment, these models demonstrate strong capabilities in content and scene understanding, enabling applications such as spoken question answering, music style analysis, and environmental sound event retrieval. They also show great promise in medical diagnosis (e.g., detection of respiratory diseases, analysis of heart sounds), voice control for smart homes, and multimedia generation and editing (Zhang et al., 2024b; Rani et al., 2017; Banerjee et al., 2024).
@@
However, the multimodal nature of ALLMs also introduces new trustworthiness challenges. First, since the models are trained on large-scale acoustic-text paired corpora, they are prone to memorizing and leaking sensitive user speech information, and are therefore vulnerable to privacy attacks such as membership inference (Tomashenko et al., 2024; Hu et al., 2022). Second, adversarial audio can exploit inaudible ultrasound or fine-grained perturbations to mislead ALLMs: early work such as DolphinAttack (Zhang et al., 2017a) and Vrifle (Li et al., 2024b) demonstrated covert manipulation of voice assistants via inaudible commands injected with ultrasonic carriers above 20 kHz (Zheng et al., 2023b; Li et al., 2023c; Ze et al., 2023); recently, AdvWave systematically proposed gradient shattering repair and two-stage optimization, achieving over 40% jailbreak success rates on various ALLMs (Kang et al., 2024). In addition, large-scale multimodal models are similarly susceptible to cross-modal instruction injection and protocol mismatching attacks, potentially leading to unauthorized content generation (Liu et al., 2023), privilege escalation (He et al., 2025), and even physical harm (Lu et al., 2024). When integrated into voice-interface agentic systems, trustworthiness challenges are amplified and become paramount (Liu et al., 2025; Yu et al., 2025). To address these risks, the community has proposed a range of safety, security, and privacy mechanisms, including SafeEar, an empirical content privacy-preserving audio deepfake detection framework (Li et al., 2024a) and active detection with post-hoc rejection (Li et al., 2023b) differentially private pre-training, segment-wise gradient compression defenses. Nevertheless, in real-time voice scenarios, these approaches still face detection latency and robustness trade-offs, highlighting the urgent need for further research.

### v4

@@ F.2 Dataset Construction Method
To construct our dataset of adversarial prompts, we followed a two-stage process: prompt generation using GPT-4o (OpenAI et al., 2024) and audio synthesis using F5-TTS (Chen et al., 2025). First, we leveraged OpenAI’s GPT-4o (the multimodal GPT-4 with voice capabilities) to automatically generate a diverse set of harmful or policy-violating user queries for each category of attack. GPT-4o was chosen for prompt generation due to its advanced language capabilities and alignment safeguards, ensuring that the produced queries are realistic and varied. We crafted a consistent prompt template to guide GPT-4o in creating these adversarial questions. This template provided the scenario context (enterprise, financial, medical, or illegal), specified the speaker’s gender (male or female) and emotional tone (angry or sad), and instructed GPT-4o to produce a user query attempting to break the system’s rules. The template was designed to simulate a malicious user’s perspective, instructing GPT-4o to create challenging queries that, while mimicking harmful intent, adhered to its content generation policies during the generation process. An example of the prompt construction template is shown below:
Using this template (with appropriate substitutions for each scenario and persona), we generated the full set of prompts. In total, we obtained 300 text prompts for the Jailbreak category and 300 for the Illegal Activity Guidance category. For the Jailbreak prompts, we generated 100 queries for each of the three scenarios (Enterprise, Financial, Medical). Within each scenario’s 100 prompts, we ensured an even coverage of the four gender–emotion combinations: 25 prompts each for Male-Angry, Male-Sad, Female-Angry, and Female-Sad. Similarly, for the Illegal Activity Guidance category, we produced 300 prompts evenly divided into 75 for each of the four combinations (male-angry, male-sad, female-angry, female-sad). This balancing ensures that our dataset controls for speaker gender and emotional tone, allowing us to analyze their effects on attack success. After generating the textual prompts, we converted each prompt into spoken audio using the F5-TTS (Chen et al., 2025) text-to-speech system. F5-TTS is a state-of-the-art TTS model capable of high-quality voice cloning with emotional expressiveness. For each prompt, we selected a voice profile that matched the intended gender and could convey the specified emotion (angry or sad). The text of the prompt was then synthesized into an audio file using F5-TTS, resulting in a spoken version of the query. This audio conversion is crucial because our attack evaluations are conducted via GPT-4o’s voice interface. Using F5-TTS to produce natural and emotive speech ensures the audio prompts closely mimic real user utterances in content and delivery.
@@ F.4 Evaluation Method and Metrics
Here, is the number of test samples, and is the GPT-4o score for the -th response. A higher DSR reflects stronger safety alignment and robustness against adversarial audio input. The safety of ALLMs is evaluated using these datasets and the safety metric detailed in Section˜F.3, with results presented in Section˜F.5.
@@ F.5 Additional Result
| Male Angry | 90.4 | 100.0 | 64.0 | 76.0 | 78.4 | 97.6 | 100.0 | 36.8 | 97.6 |
| Male Sad | 71.2 | 93.6 | 61.6 | 82.4 | 68.0 | 96.8 | 100.0 | 66.4 | 87.2 |
| Female Angry | 64.0 | 100.0 | 71.2 | 82.4 | 68.8 | 100.0 | 98.4 | 51.2 | 84.8 |
| Female Sad | 71.2 | 95.2 | 76.0 | 64.0 | 67.2 | 94.4 | 99.2 | 51.2 | 82.4 |
@@
| Male Angry | 70.4 | 68.0 | 81.6 | 80.8 | 84.0 | 89.6 | 97.6 | 48.0 | 64.0 |
| Male Sad | 72.8 | 85.6 | 78.4 | 78.4 | 90.4 | 98.4 | 98.4 | 83.2 | 65.6 |
| Female Angry | 74.4 | 94.4 | 83.2 | 76.8 | 80.0 | 96.8 | 98.4 | 56.8 | 72.0 |
| Female Sad | 80.0 | 87.2 | 79.2 | 80.8 | 90.4 | 94.4 | 98.4 | 83.2 | 65.6 |
@@
| Male Angry | 84.8 | 94.4 | 86.4 | 90.4 | 87.2 | 98.4 | 95.2 | 68.8 | 80.8 |
| Male Sad | 83.2 | 85.6 | 82.4 | 73.6 | 88.8 | 90.4 | 94.4 | 76.8 | 67.2 |
| Female Angry | 72.0 | 93.6 | 76.0 | 75.2 | 88.0 | 93.6 | 99.2 | 73.6 | 79.4 |
| Female Sad | 83.2 | 89.6 | 80.8 | 87.2 | 92.0 | 94.4 | 92.0 | 89.6 | 67.2 |
@@
| Male Angry | 84.0 | 100.0 | 98.1 | 99.2 | 99.2 | 100.0 | 100.0 | 51.2 | 73.1 |
| Male Sad | 87.7 | 96.3 | 91.7 | 96.0 | 92.3 | 99.2 | 100.0 | 85.9 | 61.9 |
| Female Angry | 72.0 | 98.7 | 89.3 | 95.5 | 91.5 | 98.7 | 99.7 | 57.3 | 71.7 |
| Female Sad | 64.8 | 97.1 | 90.7 | 94.1 | 94.9 | 100.0 | 92.8 | 92.8 | 64.0 |
@@
| Male Angry | 99.2 | 100.0 | 99.2 | 100.0 | 100.0 |
| Male Sad | 97.6 | 99.2 | 100.0 | 100.0 | 99.2 |
| Female Angry | 99.2 | 98.4 | 100.0 | 100.0 | 100.0 |
| Female Sad | 100.0 | 98.4 | 100.0 | 100.0 | 100.0 |
@@
| Male Angry | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 |
| Male Sad | 98.4 | 98.4 | 98.4 | 100.0 | 100.0 |
| Female Angry | 100.0 | 99.2 | 99.2 | 100.0 | 100.0 |
| Female Sad | 98.4 | 99.2 | 98.4 | 99.2 | 97.6 |
@@
| Male Angry | 98.4 | 98.4 | 100.0 | 100.0 | 100.0 |
| Male Sad | 96.0 | 96.8 | 98.4 | 99.2 | 99.2 |
| Female Angry | 99.2 | 98.4 | 98.4 | 100.0 | 98.4 |
| Female Sad | 96.8 | 99.2 | 97.6 | 98.4 | 100.0 |
@@
| Male Angry | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 |
| Male Sad | 100.0 | 100.0 | 100.0 | 100.0 | 100.0 |
| Female Angry | 100.0 | 100.0 | 99.7 | 99.2 | 99.7 |
| Female Sad | 99.7 | 100.0 | 100.0 | 100.0 | 99.2 |
@@ G.2 Dataset Construction Method
Direct Privacy Leakage Dataset (600 samples). For the direct privacy leakage dataset,we created six categories of sensitive personal information: bank account numbers, home addresses, phone numbers, phone passwords, driver’s license numbers, and social security numbers (Wang et al., 2023a). For each category, we first prompted Grok to generate textual data containing personal names and corresponding private information (xAI, 2025). The generated data were then further randomized to enhance diversity and reduce identifiability. Based on this preliminary dataset, Grok was instructed to generate realistic dialogue scenarios in which each type of private information might naturally appear. Subsequently, Grok produced context-appropriate conversational texts for each scenario. Finally, we selected voice samples from the Common Voice dataset and used F5-TTS to synthesize the corresponding audio files (Chen et al., 2025; Ardila et al., 2020). Each category contained 100 entries, resulting in a total of 600 test samples.
@@
Use of TTS in Privacy-Related Tasks. Within the privacy dimension of our benchmark, F5 TTS is used only for the direct privacy leakage and inference leakage (settings of conversation). For these tasks, we rely on both the systematic evaluations reported in the F5 TTS paper and our own additional checks, which demonstrate that F5 TTS produces intelligible, natural-sounding speech closely matching the reference speakers. This supports its use as a proxy for human speech in our context. In addition, we manually sampled synthesized audio for all privacy-related tasks that rely on F5 TTS. We verified that the audio faithfully matches the intended transcripts, is easy to understand, and yields clearly distinguishable voices perceived as realistic human speech. For dimensions where richer paralinguistic phenomena or real-world acoustic variability are more critical, we deliberately use real human speech or real environmental recordings instead of TTS. Taken together, the combination of manual validation and the documented realism of F5 TTS provides evidence that the constructed privacy-related datasets are an appropriate and practically useful representation of human audio interactions relevant to our evaluation.
@@ G.5 Experiment Design and Results
| Average | 80.4 | – | 70.7 | – | 77.4 | – | 90.3 | – | 78.7 | – | 74.8 | – | 23.6 | – | 29.5 | – | 56.9 | – |
Note: This table shows inference accuracy ONLY for samples where privacy was leaked. Higher values indicate greater privacy threat.
Blue arrows () indicate higher threat than average; red arrows () indicate lower threat. ‘’ indicates insufficient leaked samples ( 10).
@@ H.2 Dataset Construction Method
For Adversarial Robustness: We generate adversarial audio samples using the Projected Gradient Descent (PGD) attack method under an constraint on the waveform perturbation. The perturbation budget is set to , and 30–40 gradient steps are used for generating each adversarial example.
@@ I.2 Dataset Construction Method
Voice Cloning Spoofing (200 samples). We collected 100 samples of voice wake-up words from an existing dataset (He et al., 2022), with each sample consisting of approximately 2 seconds of wake-up word audio. These samples constitute the Voice Cloning Spoofing test set. Detailed wake-up words and their corresponding developers are listed in Table˜25. Each sample was cloned using F5-TTS (Chen et al., 2025), and we constructed paired audio clips composed of the original audio (used as the legitimate user’s enrollment input) and the cloned audio (used as the attack input attempting to bypass authentication). In addition, we designed two distinct authentication scenarios for the experiments—Awake Assistant and Open The Door. In the former, the wake-up words are used to authenticate an AI voice assistant, while in the latter, they are used to activate access to a high-security laboratory door, with an emphasis on the consequences of authentication failures. Through this experimental setup, we aim to explore how different system prompt formulations affect an ALLM’s ability to defend against voice cloning spoofing under identical audio inputs.
@@ I.4 Evaluation Method and Metrics
Our evaluation across all experiments primarily relies on the following statistical metric:
Imposter Rejection Rate (IRR), defined as the proportion of unauthorized voices that are correctly rejected by the system.
@@ I.5 Additional Result
| Ultravox | 9 | 47 | |
| Qwen2-Audio | 85 | 100 | |
| MiniCPM-o 2.6 | 73 | 86 | |
| Step-Fun | 22 | 22 | |
| Qwen2.5-omni | 0 | 0 | |
| Kimi-Audio | 8 | 41 | |
| OpenS2S | 0 | 100 | |
| Step-Audio2 | 9 | 93 | |
| Closed-source | Gemini-1.5 pro | 0 | 67 |
| GPT-4o Audio | 67 | 100 | |
| GPT-4o mini Audio | 92 | 80 | |
| Gemini-2.5 Flash | 6 | 16 | |
| Gemini-2.5-Pro | 1 | 20 |
Note: and represent the change in the defense success rate against voice cloning spoofing when comparing the ’Open the door’ scenario with stricter system prompts to the ’Awake Assistant’ scenario. Higher numbers represent better defense effectiveness.
@@
| Open-source | SALMONN | 26 | – | 7 | 8 |
| Ultravox | 95 | – | 57 | 59 | |
| Qwen2-Audio | 42 | – | 71 | 60 | |
| MiniCPM-o 2.6 | 24 | – | 43 | 56 | |
| Step-Fun | 79 | – | 97 | 98 | |
| Qwen2.5-omni | 19 | – | 64 | 36 | |
| Kimi-Audio | 79 | – | 76 | 86 | |
| OpenS2S | 97 | – | 66 | 67 | |
| Step-Audio2 | 37 | – | 15 | 20 | |
| Closed-source | Gemini-1.5 pro | 96 | 94 | 95 | 100 |
| GPT-4o Audio | 98 | 94 | 100 | 100 | |
| GPT-4o mini Audio | 100 | 91 | 100 | 100 | |
| Gemini-2.5 Flash | 97 | 96 | 93 | 81 | |
| Gemini-2.5-Pro | 95 | 100 | 97 | 100 | |
| Average | 70.3 | 95.0 | 70.1 | 69.4 | |
@@
Note: Values show imposter rejection rate (IRR) (%) with performance indicators relative to average values. indicates lower IRR than average (worse security performance); indicates higher IRR than average (better security performance). Subscripts show absolute difference from average. “–” indicates the subset of the dataset was used exclusively to evaluate model performance across different modalities (audio and text). The evaluation was restricted to closed-source models to streamline the analysis. “bg” = with background audio, “no bg” = without background audio.
@@
Voice Cloning Spoofing. In Table˜26, we analyze the experimental results of all open-source and closed-source models under two scenarios with different levels of text prompt flexibility. It can be observed that most models perform better in the “Open The Door” scenario than in the “Awake Assistant” scenario. with a significant increase in the defense success rate against voice cloning spoofing. This is particularly evident for Ultravox, Gemini-1.5 Pro, and Gemini-2.5 Pro. This indicates that even in Audio-based Large Language Models (ALLMs) where audio is the primary input, the accuracy of text prompts still plays a significant role. Furthermore, this has implications for the downstream applications of ALLMs: for scenarios involving security, authentication, etc., designing a strict and precise prompt may lead to a considerable improvement in model performance.
@@
Identity Verification Bypass. From the Table˜27, it can be observed that closed-source models are harder to deceive compared to open-source models. Among them, GPT-4o mini Audio performs the best, with a IRR(Imposter Rejection Rate) of only 100%. Among all closed-source models, Qwen2.5-omni performs the worst, with a IRR as low as 19%. These results indicate that even without providing complete or explicit authentication information, voice models still have a high probability of passing identity verification, which poses a significant security risk. In the Table˜27, we also investigated the IRR metric in pure text mode, which is labeled as the “Text” column. This represents using the corresponding text-based model of the audio model to perform inference on the text version of the identity verification bypass dataset. It can be observed that, in general, the IRR is lower in text mode compared to audio mode. This suggests that the additional paralinguistic information present in the speech modality, such as emotional cues or prosodic features, may contribute positively to the authentication performance of the model.
@@
Hybrid Spoofing. In this configuration, we simulate social engineering attacks combined with background audio that may occur in real authentication scenarios, aiming to study the impact of background sounds on the verification outcome. The experimental results show that the influence of added background audio on model performance does not follow an obvious pattern. For instance, Qwen2-Audio’s IRR increases by 11%, whereas Gemini-1.5 Pro’s IRR decreases instead.
@@
With the rapid increase in parameter and data scales, text-only large language models (LLMs) have achieved groundbreaking progress in language understanding and generation, as exemplified by models such as GPT‑4 and the Gemini series (Achiam et al., 2023; Team et al., 2023). Building on this, researchers explored cross-modal alignment by integrating visual information into unified representation spaces. This led to early models like CLIP (Radford et al., 2021) and Flamingo (Alayrac et al., 2022), and later, models such as GPT-4V and Gemini capable of processing high-resolution images and long contexts. Recently, ALLMs have further expanded the input modalities by incorporating temporal acoustic features (such as Mel-spectrograms, log-power spectra, or variable-length waveforms) for joint modeling with semantic tokens (Yang et al., 2024b). In contrast to the visual modality, audio signals exhibit high dynamic range and transient variations in both time and frequency domains. Consequently, most ALLMs adopt separate time-frequency encoders or discretizing acoustic tokenizers to capture rich attributes such as timbre, rhythm, and scene (Hu et al., 2024; Du et al., 2023). Representative models include Qwen2‑Audio with its pipeline-style natural language prompt pre-training (Chu et al., 2024), SALMONN with a unified "auditory-language-music" framework (Tang et al., 2024), and WavLLM with a dual-encoder plus Prompt-LoRA adaptation mechanism (Hu et al., 2024). After cross-modal alignment, these models demonstrate strong capabilities in content and scene understanding, enabling applications such as spoken question answering, music style analysis, and environmental sound event retrieval. They also show great promise in medical diagnosis (e.g., detection of respiratory diseases, analysis of heart sounds), voice control for smart homes, and multimedia generation and editing (Zhang et al., 2025; Rani et al., 2017; Banerjee et al., 2024).
@@
However, the multimodal nature of ALLMs also introduces new trustworthiness challenges. First, since the models are trained on large-scale acoustic-text paired corpora, they are prone to memorizing and leaking sensitive user speech information, and are therefore vulnerable to privacy attacks such as membership inference (Tomashenko et al., 2024; Hu et al., 2022). Second, adversarial audio can exploit inaudible ultrasound or fine-grained perturbations to mislead ALLMs: early work such as DolphinAttack (Zhang et al., 2017a) and Vrifle (Li et al., 2024b) demonstrated covert manipulation of voice assistants via inaudible commands injected with ultrasonic carriers above 20 kHz (Zheng et al., 2023b; Li et al., 2023c; Ze et al., 2023); recently, AdvWave systematically proposed gradient shattering repair and two-stage optimization, achieving over 40% jailbreak success rates on various ALLMs (Kang et al., 2025). In addition, large-scale multimodal models are similarly susceptible to cross-modal instruction injection and protocol mismatching attacks, potentially leading to unauthorized content generation (Liu et al., 2023), privilege escalation (He et al., 2025), and even physical harm (Lu et al., 2024). When integrated into voice-interface agentic systems, trustworthiness challenges are amplified and become paramount (Liu et al., 2025; Yu et al., 2025). To address these risks, the community has proposed a range of safety, security, and privacy mechanisms, including SafeEar, an empirical content privacy-preserving audio deepfake detection framework (Li et al., 2024a) and active detection with post-hoc rejection (Li et al., 2023b) differentially private pre-training, segment-wise gradient compression defenses. Nevertheless, in real-time voice scenarios, these approaches still face detection latency and robustness trade-offs, highlighting the urgent need for further research.